The sum of two numbers is 118.
  
If three times

the smaller number is subtracted from the larger​ number, the result is 18.
  
Find the two numbers.

Found 2 solutions by addingup, solver91311:
Answer by addingup(3677)

x+y =118
x-3y = 18 multiply all by -1 and add to the other equation:
````````````````````````````````````````````````````
x+y = 118
+
-x+3y = -18
4y = 100
y = 25
Now let's find x:
x-3y = 18
x-3(25) = 18
x-75 = 18
x = 93
Check
x+y = 118
93+25 = 118 Correct.
John

Answer by solver91311(24713)







Solve for and . Hint: Multiply the 2nd equation by -1 and then use elimination.

John

My calculator said it, I believe it, that settles it